10. Consider the following summary statistics. Using the boxplot outlier test, potential outliers would be any value above what?: *

(A) 156

(B) 183

(C) 255

(D) 282

Questions 11-12: Consider the simple linear regression model y = a + bx. Now select a sample size of 10 with the following information:

Σ x Σ = 56 y Σ ( = 67 x – x )( y – y Σ ( ) = 780 x – x ) 2 = 1250

11. What is the estimated value of b?: *

(A) 0.50

(B) 0.62

(C) 0.72

(D) 1.00

12. What is the estimated value of a?: *

(A) 1.8

(B) 2.9

(C) 3.2

(D) 4.5
Answer #1

The solution of the system can be presented as:

1586433967361_blob.png= 56 / 10 = 5.6

1586433967210_blob.png = 67/ 10 = 6.7

1586433967269_blob.png

1586433967422_blob.png

1586433967179_blob.png

b = \frac{Sxy}{Sxx}

given-

Σ x = 56

Σ y = 67

Σ (x – x )( y – y) = 780

Σ (x – x ) 2 = 1250

11)

b = \frac{Sxy}{Sxx}

b = Σ (x – x )( y – y) / Σ (x – x ) 2

b = 780 / 1250 = 0.62

12)

1586433967179_blob.png

a = 6.7 - (0.62 * 5.6)

a = 6.7 - 3.472

a = 3.2

please like)